A solution is 0.1MBa(No3) 2 & 0.1Msr(No3)2. If solid Na2cro4 is added to the solution. What is [Ba] when srcro4 begins to precipitate?? Ksp of Bacro4 is 1.2×10^(-10) , Ksp of srcro4 is 3.5×10^(-5).

A solution is  0.1MBa(No3) 2 & 0.1Msr(No3)2. If solid Na2cro4 is added to the solution. What is [Ba] when srcro4 begins to precipitate?? Ksp of  Bacro4 is 1.2×10^(-10) , Ksp of srcro4 is 3.5×10^(-5).

Grade:11

1 Answers

B Gowreeswaramma
5 years ago
[Ba] = ksp / [cro4],   [cro4=ksp/[ Sr]  i. e., 3.5 *10^(-5) /0.1    , =3.5 *0.0001 so Ba =3.7*10^(-7)approximatelya